OpenStudy (anonymous):

If z=cos(x), find the second derivative with respect to x . i.e. d^{2}y/dx^{2}

OpenStudy (anonymous):

do you mean: \[\frac{d^2z}{dx^2}\] (z instead of y)

OpenStudy (anonymous):

No, this is the exact question: If z=sin(x) and if y is a function of x, show that d^{2}y/dx^{2}=cos^{2}xd^{y}/dz^{2}-sin(x)dy/dz

OpenStudy (anonymous):

I first found dy/dx using the chain rule. dy/dx=dy/dz.dz/dx=cos(x)dy/dz I proceeded to find the second derivate w.r.t x but it does not match that given.